OpenStudy (anonymous):

find the quotient -5y^3+25y/-5y

Directrix (directrix):

Take the numerator: -5y^3+25y and factor out the greatest common monomial factor. -5y^3+25y = -5y*( y² - ?) Post what you get. Then, it will be time to divide the factored numerator by the denominator: (-5y) @CollinsJennifer

OpenStudy (anonymous):

I still don't understand

Directrix (directrix):

Factor this: -5y^3+25y = -5y*( y² - ?) -5y times what is equal to 25y ?

OpenStudy (anonymous):

-5

Directrix (directrix):

So, -5y^3+25y = -5y*( y² - 5)

Directrix (directrix):

 [-5y*( y² - 5]) divided by (-5y)